OpenStudy (kpop4life123):

PLEASE HELP ME!!!!! Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 4−x and y = 2x + 3 intersect are the solutions of the equation 4−x = 2x + 3. (4 points) Part B: Make tables to find the solution to 4−x = 2x + 3. Take the integer values of x between −3 and 3. (4 points)

OpenStudy (tarnett1):

Part A: In order to distinguish the two line, make the Y in (formula of) the first line a Y1 and the other Y2. Y1= -x +2 Y2= 4x +3 Now when the two lines intersect the values for the Y's at that point must be equal. So -x +2 = Y1 = Y2 = 4x +3 Part B: What? Make tables? Doesn't you teacher know algebra? -x +2 = 4x +3 Isolate the x's 2-3 = 4x +x -1 = 5x x = -1/5 = - 0.2 ==> Y1 = Y2= -(-0.2) + 2 = 2.2 So the lines should intersect at the point (-0.2, 2.2) Part C: To do this (easily) put both equations in slope-(y)intercept form. y = mx +b In this form m= the slope (rise/run) and b= the y-intercept = the point on the y-axis where the line crosses the y-axis. So for Y1= -x +2 , put a little mark on the vertical (Y) axis at +2, now from that point, count 2 (unit spaces) to the right, then from there count down 2 spaces and put a little mark. Now connecting those 2 points draws the first line. For Y2= 4x+3, find 3 units up on the y-axis (mark), now from there move to the right 3 units, From there move up 12 spaces and put a mark. I used 3 and 12 because m=4=12/3. Now connect the two marks to draw the second line. Again, the lines should intersect at the point (-0.2, 2.2)

OpenStudy (kpop4life123):

is this your own work?

OpenStudy (tarnett1):

why?

OpenStudy (kpop4life123):

cuz i can't use the ones on the other sites

OpenStudy (tarnett1):

what do you mean

OpenStudy (kpop4life123):

reasons

OpenStudy (kpop4life123):

can someone please help???

OpenStudy (cristic):

may you just try draw the y = -x +2 and y = 4x +3

OpenStudy (tarnett1):

well it wasnt me that posted that it was my sister she gets o my account and i dont know if she answered it her self

OpenStudy (cristic):



OpenStudy (cristic):

just try